OpenStudy (anonymous):

Solve x please show steps 0.00632+0.00632x=0.01-x

OpenStudy (anonymous):

0.00632+0.00632x=0.01-x 1.00632x = 0.01 - 0.00632 x = (0.01-0.00632) / 1.00632 = 0.00365

OpenStudy (anonymous):

How'd you get 1.00632x?? @antoni7

OpenStudy (anonymous):